OpenStudy (anonymous):

if two numbers x and y are chosen at random from the set of first 30 natural numbers. the chance that x^2-y^2 is divisible by 3 is

OpenStudy (anonymous):

Cha-Ching, awards in the bank.

OpenStudy (anonymous):

x^2-y^2= (x+y)(x-y) so we need to check what is the chance that x+y or x-y is divisible of 3 if x is divisible of 3 (1/3 chance) than there is 1/3 chance that x+y will be divisible and in that case both are (so x-y) too if x is not divisible of 3 (2/3)than if y is not divisible of 3 (2/3) chance than either x+y or x-y will be factor of 3. If y is divisible of 3 than neither x+y and x-y will be divisible. So the answer is 1/3*1/3+2/3*2/3=5/9 I am completely not sure if this is correct :-) But I thought about it for some time.

OpenStudy (anonymous):

no the answer is 47/87

OpenStudy (anonymous):

can you see where my logic failed?

OpenStudy (anonymous):

"if x is divisible of 3 (1/3 chance) than there is 1/3 chance that x+y will be divisible and in that case both are (so x-y) too" why do you say 1/3 chance?

OpenStudy (anonymous):

that x+3 and x-y are divisible of 3

OpenStudy (anonymous):

x+3?

OpenStudy (anonymous):

x+y

OpenStudy (anonymous):

how does dividing x by 3 has one-third chance?? m not gettin it..

OpenStudy (anonymous):

1/3*9/29+2/3*19/29= 47/87

OpenStudy (anonymous):

my logic was good but I did not see that it is without replacement.

OpenStudy (anonymous):

1/3 chance that x is divisible of 3 than x+y or x-y only divisible of 3 if y is divisible of 3, that is 9/29 2/3 chance that x is not divisible of 3 than if y is not divisible of 3 than either x+y or x-y will be. and that is 19/29